# Ground States of ${\text{Class}\;\mathcal{S}}$ Theory on ADE Singularities and dual Chern-Simons theory

### v1

@@ 3.2.1 Comparison with $SO(8)_{2}$ Chern-Simons theory
The Chern-Simons theory to which we need to compare the results of §3.2, under the VW duality, has an SO$(8)$ gauge group (dual to Dic2 as given by the McKay correspondence in Table 1) and level $2$ [since our initial gauge group is SU$(2)$]. The 5 possible holonomies labeled by $a=1,\dots,5$ correspond to points lying on the lattice (2), which in terms of SO$(8)$ Dynkin labels (i.e., the coefficients in an expansion in the fundamental weights) are given by:
@@
We expect states with energy $2E_{0}$ to correspond to operators with dimension $\Delta=0$, and we will not discuss operators with dimension $\Delta>0$ in this paper, but we do need to discuss the operators with dimension $\Delta<0$ that arise from gluing 3-punctured spheres with $(555)$ triplets and energy $E_{3}=E_{0}-\frac{1}{2}$.
@@ 5 Generalization and conclusion
In this work, we put forward a conjecture about the action of the duality group on dimension $\Delta=0$ operators of a ${\text{Class}\;\mathcal{S}}$ theory formulated on $\mathbb{R}^{4}/\Gamma$, where the operators are at the singularity. The conjecture, motivated by Vafa and Witten’s results for ${\mathcal{N}}=4$ SYM Vafa:1994tf is that it is equivalent to the action of the mapping class group of the Riemann surface $\Sigma$ (underlying Gaoitto’s construction of the ${\text{Class}\;\mathcal{S}}$ theory) on the Hilbert space of Chern-Simons theory on $\Sigma$ with gauge group determined by $\Gamma$, through the McKay correspondence, and the level determined by the gauge group of the ${\text{Class}\;\mathcal{S}}$ theory. We provided some modest supporting evidence by matching the number of ground states of the relevant Chern-Simons theory with the number of $\Delta=0$ operators at the singularity, for $\Gamma=\mathbb{Z}_{k}$ and $\Gamma=\text{Dic}_{2}$, corresponding to Chern-Simons gauge groups SU$(k)$ and SO$(8)$. The counting for the abelian $\mathbb{Z}_{k}$ proceeded in the same way as previously computed in the literature in the context of the 4d/2d and 3d/3d correspondence Benini:2011nc ; Alday:2013rs ; Gukov:2015sna , with the difference that our Chern-Simons theory parameters are connected by level-rank duality to a compact form of the noncompact Chern-Simons theory that appears in the 3d/3d correspondence, but the calculation is otherwise identical. It relies on the nontrivial Casimir energies that arise for some combinations of flat connections when the ${\text{Class}\;\mathcal{S}}$ theory is formulated on $S^{3}/\Gamma$. (For $N=4$ SYM, such a Casimir energy does not produce constraints on holonomies, as demonstrated explicitly in Ju:2023ssy .) A novel aspect of our results, to the best of our knowledge, is the extension of the calculation to a nonabelian $\Gamma=\text{Dic}_{2}$ where the $\Delta=0$ operators again match with Chern-Simons states, thanks to restrictions due to nontrivial Casimir energies. For the nonabelian $\Gamma$, we also found a state with $\Delta<0$, which is another novelty. It would be interesting to explore correlation functions of these operators, say, on $T^{4}/\text{Dic}_{2}$, where the action of $\text{Dic}_{2}$ is defined so as to produce two singularities that look locally as $\mathbb{R}^{4}/\text{Dic}_{2}$, as well as additional singularities of type $A_{1},A_{3}$. Extensions to other nonabelian $\Gamma$’s, i.e., corresponding to $E_{6},E_{7},E_{8}$, or $\text{Dic}_{k}$ with $k>2$, will be explored elsewhere. Extensions to ${\text{Class}\;\mathcal{S}}$ theories derived from D-type or E-type $(2,0)$-theory are also left for future exploration.
@@ Appendix B Details of $S^{3}/\text{Dic}_{2}$ index calculation
In this section, we give a detailed computation of the superconformal single letter index on $S^{3}/\text{Dic}_{2}$. The idea behind the computation is discussed in §3.2. A quick summary is that, for a BPS operator to contribute to the single letter index, it must satisfy two constraints as in equations (42) and (43). The contribution of an operator to the index is given by the Boltzmann weight $t^{2(\Delta+j_{2})}$, where for each field its charges $(\Delta,j_{2})$ are given in Table 3. In this appendix, we will explicitly solve for the constraints and sum over the appropriate Boltzmann weights in order to find the index in the case of the $\text{Dic}_{2}$ group. We recall that are five flat connections Table 4 which we label by $j\in\{1,2,3,4,5\}$. The first four are abelian, with $j=1$ labeling a trivial one, and $j=5$ is nonabelian. The nontrivial abelian flat connections are $j=2,3,4$. Counting triplets of holonomies, we will find that in total, there are 125 triplets, given by the following groupings:
@@
  • $(115)$: 3 triplets involving two trivial holonomies and one nonabelian holonomy. There are 3 holonomy triplets of this type due to permutation.
@@
$r\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi$ $=e^{i(n_{1}-n_{2})/2}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi,$
$s\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi$ $=e^{\pi in_{2}}\partial^{n_{2}}_{++}\partial^{n_{1}}_{-+}\Phi.$
@@
$r\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\lambda_{\mu}$ $=e^{i(n_{1}-n_{2}+\mu)/2}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\lambda_{\mu},$
$s\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\lambda_{\mu}$ $=(-1)^{(\mu-1)/2}e^{\pi in_{2}}\partial^{n_{2}}_{++}\partial^{n_{1}}_{-+}\lambda_{-\mu},$
@@
In §3.2 We discussed the effect of the gauge transformation experienced by the fields along a nontrivial $1$-cycle of $S^{3}/\Gamma$. See equations (50) and (51) for the vector multiplet and equation (58) for the hypermultiplet. Therefore, we now have all the ingredients we need to compute the single letter indices. Since the three vector multiplets do not couple to each other, we can compute the vector multiplet contributions to the index for each individual holonomy in the holonomy triplet. We begin by computing the vector multiplet contribution to the Index for all holonomies (1-5).
B.1 Contribution from vector multiplet $I^{vec}$
@@
First, we note that the vector multiplet contributions for holonomies 1-4 is the same
$I^{vec}_{1}=I^{vec}_{2}=I^{vec}_{3}=I^{vec}_{4}.$
@@
This is because these holonomies lead to the same adjoint actions. We first compute the contribution from the BPS fields $\Phi$ that have spin $j_{1}=0$. As discussed in §3.2, we need a linear combination of the descendants to satisfy the constraints:
$\sum c^{n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi.$
The algebraic structure of the Dicyclic group leads to a constraint between $r$ and $s$, which results in the equations:
$\sum c^{n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi$ $=\sum e^{i\pi/2(n_{1}-n_{2})}c^{n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi,$
$\sum c^{n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi$ $=\sum e^{i\pi n_{2}}c^{n_{1}n_{2}}\partial^{n_{2}}_{++}\partial^{n_{1}}_{-+}\Phi.$
The first equation implies that, for $c^{n_{1}n_{2}}$ to be nonvanishing, we require
$n_{1}-n_{2}=0\mod{4},$
and the second constraint shows that we only need to sum over $n_{1}\geq n_{2}$. However, for $n_{1}=n_{2}$, we need to sum over $n_{1}=0\mod{2}$ only, since the coefficient vanishes otherwise. Therefore, the constrained sum:
$F=3\sum t^{3n_{1}}t^{3n_{2}},$
becomes
$F=3\frac{1-t^{6}+t^{12}}{(1-t^{6})(1-t^{12})},$
where the prefactor 3 below comes from the 3 SU$(2)$ adjoint degrees of freedom and contribution of each letter is read from Table 3. Hence, the contribution of the spin projection $j_{1}=0$ fields in the vector multiplet to the single letter index for holonomy 1-4 is:
$F(t^{2}-t^{4}+2t^{6}).$ (64)
@@
For the case of fields $\lambda_{\mu}$ with a spin projection $j_{1}=\pm 1/2$, a generic operator now has the following form:
$\sum c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\lambda_{\mu}.$
@@
$\sum c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\lambda_{\mu}$ $=\sum e^{\pi i/2(n_{1}-n_{2}+\mu)}c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\lambda_{\mu},$
$\sum c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\lambda_{\mu}$ $=\sum(-1)^{(\mu-1)/2}e^{\pi in_{2}}c^{\mu n_{1}n_{2}}\partial^{n_{2}}_{++}\partial^{n_{1}}_{-+}\lambda_{-\mu}.$
The first constraint shows that when $\mu=1$, we have
$n_{1}-n_{2}=3\mod{4},$
and the second constraint shows that we need to sum over all $n_{1}$ and $n_{2}$ subject to the above equation. We do not have to do this again for the $\mu=-1$ case, since $c^{-1,n_{1}n_{2}}$ is completely determined from $c^{1,n_{1}n_{2}}$ from the second constraint. Therefore, the sum becomes
$F^{\prime}=3\sum t^{3n_{1}}t^{3n_{2}}=3\frac{t^{3}+t^{9}}{(1-t^{6})(1-t^{12})},$
and the contribution of the $\lambda_{\pm}$ fields to the index is
$-F^{\prime}t^{3}.$ (65)
Adding (64) and (65), we find that the vector multiplet contribution to the single letter index for holonomy 1-4 is
$I^{vec}_{1}$ $=F(t^{2}-t^{4}+2t^{6})-F^{\prime}t^{3}$
$=\frac{3(1-t^{6}+t^{12})(t^{2}-t^{4}+2t^{6})-3(t^{3}+t^{9})t^{3}}{(1-t^{6})(1-t^{12})}.$
Next, we compute $I^{vec}_{5}$, the vector multiplet single letter index for the 5th holonomy. This case is harder because the 5th holonomy is non-abelian. As discussed in §3.2, under $g_{5}(r)$ and $g_{5}(s)$, the adjoint fields $\Phi_{p}$, with $p=\{0,-1,1\}$ transform as
$g_{5}(r)(\Phi_{0},\Phi_{1},\Phi_{-1})g^{-1}_{5}(r)$ $=(\Phi_{0},-\Phi_{1},-\Phi_{-1}),$
$g_{5}(s)(\Phi_{0},\Phi_{1},\Phi_{-1})g^{-1}_{5}(s)$ $=(-\Phi_{0},-\Phi_{-1},-\Phi_{1}).$
As before, we do the computation for fields with spin projection $j_{1}=0$ and $j_{1}=\pm 1/2$ separately. A generic operator that corresponds to a spin $j_{1}=0$ field has the form:
$\sum c^{pn_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{p}.$
@@
$\sum e^{i\pi/2(n_{1}-n_{2})}c^{pn_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{p}$ $=\sum(-1)^{p}c^{pn_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{p},$
$\sum e^{i\pi n_{2}}c^{pn_{1}n_{2}}\partial^{n_{2}}_{++}\partial^{n_{1}}_{-+}\Phi_{p}$ $=-\sum c^{pn_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{-p}.$
@@
$e^{i\pi/2(n_{1}-n_{2})}c^{pn_{1}n_{2}}$ $=(-1)^{p}c^{pn_{1}n_{2}},$
$e^{i\pi n_{1}}c^{-p,n_{2}n_{1}}$ $=-c^{pn_{1}n_{2}}.$
There are two cases to consider:
Case 1: $p=0$. For this case, the first constraint shows that
$n_{1}-n_{2}=0\mod{4},$
and the second constraint instructs us to sum over only $n_{1}\geq n_{2}$, but for $n_{1}=n_{2}$ we need only sum over $n_{1}=1\mod{2}$. Therefore, the sum becomes
$F^{0}=\frac{t^{6}}{(1-t^{6})(1-t^{12})}.$ (66)
Case 2: $p=1$. For this case, the first constraint shows that
$n_{1}-n_{2}=2\mod{4},$
and the second constraint tells us to sum over all such $n_{1}$ and $n_{2}$ obeying the above equation. In addition, it implies that we do not have to consider the $p=-1$ case, since it is already included. Therefore, the sum is
$F^{+}=\frac{2t^{6}}{(1-t^{6})(1-t^{12})}=2F^{0}.$ (67)

### v2

@@ 3.2.1 Trinion contribution to the Casimir energy on $S^{3}/\text{Dic}_{2}$
3.2.2 Combining the vertex contributions for $\Sigma$
We decompose the Riemann surface $\Sigma$ into trinions, label each edge by a flat connection $1,\dots,5$, and add the contributions of the vertices according to the list at the end of §3.2.1. For example, if the edges in Figure 1 are $a=1$ and $b=c=2$ we need twice the contribution of the (122) triplet, and the Casimir energy is $2E_{0}$, while if $a=b=1$ and $c=2$ we get $42E_{0}+$ 4, and if $a=b=c=5$, we get $12E_{0}-$ 1.
In general, for a Riemann surface of genus $g>1$, if all the vertices ($22g-$ 2 in number) are labeled by triplets from the first item of the list in §3.2.1, the resulting Casimir energy will be $E=(2g-2)E_{0}$. This is also the Casimir energy for the trivial flat connection (i.e., all edges labeled by $1$), and we expect these states to correspond to dimension $\Delta=0$ operators on $\mathbb{R}^{4}/\Gamma$, where $(2g-2)E_{0}$ is accounted for by the conformal anomaly of the state-operator transformation. We will argue in §3.2.3 that such states precisely agree with the conjectured Chern-Simons dual system.
For genus $g\leq 3$, all other combinations of labels on the edges lead to states with Casimir energy $E>(2g-2)E_{0}$, except when all edges are labeled by $5$, in which case the energy is $E=(2g-2)E_{0}+1-g$, which we expect to lead to $\Delta=1-g<0$. This does not necessarily contradict unitarity, as we will discuss in §3.2.4, but these states are not captured by the proposed Chern-Simons system.
For $g\geq 4$, we can also combine vertices with Casimir energy $E_{0}+2$ and vertices with Casimir energy $E_{0}-\frac{1}{2}$ to get a total energy of $E=(2g-2)E_{0}$, just like the trivial flat connection. These states also do not fit with the proposed Chern-Simons dual, as we explain in §3.2.3. We expect that under the action of the mapping class group of $\Sigma$, those states do not mix with those that are constructed purely from combinations of the triplets from the first item of the list in §3.2.1:
$(111),(122),(133),(144),(234),(155),(255),(355),(455).$
@@ 3.2.3 Comparison with $SO(8)_{2}$ Chern-Simons theory
The Chern-Simons theory to which we need to compare the results of §3.2.1, under the VW duality, has an SO$(8)$ gauge group (dual to Dic2 as given by the McKay correspondence in Table 1) and level $2$ [since our initial gauge group is SU$(2)$]. The 5 possible holonomies labeled by $a=1,\dots,5$ correspond to points lying on the lattice (2), which in terms of SO$(8)$ Dynkin labels (i.e., the coefficients in an expansion in the fundamental weights) are given by:
@@
We expect states with energy $2E_{0}$ to correspond to operators with dimension $\Delta=0$, and we will not discuss operators with dimension $\Delta>0$ in this paper. We leave the question of the action of the mapping class group of $\Sigma$ on the other states discussed in §3.2.2, including those with $\Delta<0$, as an open problem.
@@ 5 Generalization and conclusion
In this work, we put forward a conjecture about the action of the duality group on dimension $\Delta=0$ operators of a ${\text{Class}\;\mathcal{S}}$ theory formulated on $\mathbb{R}^{4}/\Gamma$, where the operators are at the singularity. The conjecture, motivated by Vafa and Witten’s results for ${\mathcal{N}}=4$ SYM Vafa:1994tf is that it is equivalent to the action of the mapping class group of the Riemann surface $\Sigma$ (underlying Gaiotto’s construction of the ${\text{Class}\;\mathcal{S}}$ theory) on the Hilbert space of Chern-Simons theory on $\Sigma$ with gauge group determined by $\Gamma$, through the McKay correspondence, and the level determined by the gauge group of the ${\text{Class}\;\mathcal{S}}$ theory. We provided some modest supporting evidence by matching the number of ground states of the relevant Chern-Simons theory with the number of $\Delta=0$ operators at the singularity, for $\Gamma=\mathbb{Z}_{k}$ and $\Gamma=\text{Dic}_{2}$, corresponding to Chern-Simons gauge groups SU$(k)$ and SO$(8)$. The counting for the abelian $\mathbb{Z}_{k}$ proceeded in the same way as previously computed in the literature in the context of the 4d/2d and 3d/3d correspondence Benini:2011nc ; Alday:2013rs ; Gukov:2015sna , with the difference that our Chern-Simons theory parameters are connected by level-rank duality to a compact form of the noncompact Chern-Simons theory that appears in the 3d/3d correspondence, but the calculation is otherwise identical. It relies on the nontrivial Casimir energies that arise for some combinations of flat connections when the ${\text{Class}\;\mathcal{S}}$ theory is formulated on $S^{3}/\Gamma$. (For ${\mathcal{N}}=4$ SYM, such a Casimir energy does not produce constraints on holonomies, as demonstrated explicitly in Ju:2023ssy .) A novel aspect of our results, to the best of our knowledge, is the extension of the calculation to a nonabelian $\Gamma=\text{Dic}_{2}$ where the $\Delta=0$ operators again match with Chern-Simons states, thanks to restrictions due to nontrivial Casimir energies. For the nonabelian $\Gamma$, we also found a state with $\Delta<0$, which is another novelty. It would be interesting to explore correlation functions of these operators, say, on $T^{4}/\text{Dic}_{2}$, where the action of $\text{Dic}_{2}$ is defined so as to produce two singularities that look locally as $\mathbb{R}^{4}/\text{Dic}_{2}$, as well as additional singularities of type $A_{1},A_{3}$. Extensions to other nonabelian $\Gamma$’s, i.e., corresponding to $E_{6},E_{7},E_{8}$, or $\text{Dic}_{k}$ with $k>2$, will be explored elsewhere. Extensions to ${\text{Class}\;\mathcal{S}}$ theories derived from D-type or E-type $(2,0)$-theory are also left for future exploration.
@@ Appendix B Details of $S^{3}/\text{Dic}_{2}$ index calculation
In this section, we give a detailed computation of the superconformal single letter index on $S^{3}/\text{Dic}_{2}$. The idea behind the computation is discussed in §3.2. A quick summary is that, for a BPS operator to contribute to the single letter index, it must satisfy two constraints as in equations (42) and (43). The contribution of an operator to the index is given by the Boltzmann weight $t^{2(\Delta+j_{2})}$, where for each field its charges $(\Delta,j_{2})$ are given in Table 3. In this appendix, we will explicitly solve for the constraints and sum over the appropriate Boltzmann weights in order to find the index in the case of the $\text{Dic}_{2}$ group. We recall that there are five distinct flat connections (see Table 4) which we label by $j\in\{1,2,3,4,5\}$. The first four are abelian, with $j=1$ labeling the trivial connection, and $j=5$ is (the only) nonabelian connection. The nontrivial abelian flat connections are $j=2,3,4$. Counting triplets of holonomies, we will find that in total, there are 125 triplets, given by the following groupings:
@@
  • $(115)$: 3 triplets involving two trivial holonomies and one nonabelian holonomy. There are 3 holonomy triplets of this type, counting permutations as distinct triplets.
@@
$r\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi$ $=e^{i\pi(n_{1}-n_{2})/2}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi,$
$s\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi$ $=(-1)^{n_{2}}\partial^{n_{2}}_{+\dot{+}}\partial^{n_{1}}_{-\dot{+}}\Phi.$
@@
$r\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\lambda_{\mu}$ $=e^{i\pi(n_{1}-n_{2}+\mu)/2}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\lambda_{\mu},$
$s\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\lambda_{\mu}$ $=(-1)^{(\mu-1)/2}e^{\pi in_{2}}\partial^{n_{2}}_{+\dot{+}}\partial^{n_{1}}_{-\dot{+}}\lambda_{-\mu},$
@@
In §3.2 We discussed the effect of the gauge transformation experienced by the fields along a nontrivial $1$-cycle of $S^{3}/\Gamma$. See equations (50) and (51) for the vector multiplet and equation (58) for the hypermultiplet. Therefore, we now have all the ingredients we need to compute the single letter indices. For the contribution of the vector multiplets, we simply compute the contribution of each multiplet separately, and then multiply the results for the three vector multiplets corresponding to the three $SU(2)$ gauge group factors. The contribution of each vector multiplet, separately, depends only on one of the holonomies (labeled by $j=1,\dots,5$) of the holonomy triplet, and we begin by computing that contribution.
B.1 Contribution of a single vector multiplet $I^{vec}$
@@
First, we note that the vector multiplet contributions for holonomies $1-4$ are identical,
$I^{vec}_{1}=I^{vec}_{2}=I^{vec}_{3}=I^{vec}_{4},$
because when acting on fields in the adjoint representation of the gauge group $SU(2)$, the representations of $\text{Dic}_{2}$ defined by each of the holonomies $1-4$ are all trivial, and there is no distinction between the four holonomies $1-4$ (see Table 4).
@@
We first compute the contribution from the BPS fields $\Phi$ that have spin $j_{1}=0$. As discussed in §3.2, we need to compute the dimension of the vector space of linear combinations of the descendants, of the form
$\sum c^{n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi,$ (64)
with constant coefficients $c^{n_{1}n_{2}}$, that satisfy two constraints which follow from the action of the generators $r,s$ of the Dicyclic group:
$\sum c^{n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi$ $=\sum e^{i\pi/2(n_{1}-n_{2})}c^{n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi,$
$\sum c^{n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi$ $=\sum e^{i\pi n_{2}}c^{n_{1}n_{2}}\partial^{n_{2}}_{+\dot{+}}\partial^{n_{1}}_{-\dot{+}}\Phi.$
The first equation implies that $c^{n_{1}n_{2}}$ must vanish unless
$n_{1}-n_{2}\equiv 0\pmod{4},$
and the second constraint shows that $c^{n_{1}n_{2}}$ with $n_{1}<n_{2}$ is determined by $c^{n_{1}n_{2}}$ with $n_{1}>n_{2}$ and as for the diagonal terms, $c^{n_{1}n_{1}}$ vanish unless $n_{1}$ is even.
Using Table 3, we find that ++˙$\partial^{n_{2}}_{+\dot{+}}\partial^{n_{1}}_{-\dot{+}}$ contributes $t^{3n_{1}+3n_{2}}y^{n_{2}-n_{1}}$ to the single letter index. Let $I_{\Phi}(t)$ be the contribution of the (as yet unspecified) field $\Phi$ to the single-letter index. Since $\Phi$ is assumed to be in the adjoint representation of $SU(2)$, which is $3$-dimensional, we include an extra factor of $3$ and find the contribution of words of the form (64) to the single-letter index to be, after setting $y=1$,
$F=3\left\{\sum_{\tiny\begin{array}[]{c}n_{1}>n_{2}\geq 0\\ n_{1}\equiv n_{2}\!\!\!\!\pmod{4}\\ \end{array}}t^{3n_{1}+3n_{2}}+\sum_{\tiny n_{1}\equiv 0\!\!\!\!\pmod{2}}t^{6n_{1}}\right\}I_{\Phi}(t)=\frac{3(1-t^{6}+t^{12})}{(1-t^{6})(1-t^{12})}I_{\Phi}(t)\,.$ (65)
Referring to Table 3, the possible fields to substitute for $\Phi$ are those with $j_{1}=0$, which are $\phi$, $\overline{\lambda}_{1\dot{+}}$, and $\overline{F}_{\dot{+}\dot{+}}$. Setting the fugacity $y=1$, we find $I_{\Phi}(t)=t^{2}-t^{4}+t^{6}$, and therefore, the contribution of the spin projection $j_{1}=0$ fields in the vector multiplet to the single letter index for holonomies $1-4$ is given by
$\frac{3(1-t^{6}+t^{12})(t^{2}-t^{4}+t^{6})}{(1-t^{6})(1-t^{12})}\,.$ (66)
@@
For the case of fields $\lambda_{\mu}$ ($\mu=\pm 1$) with spin projections $j_{1}=\mu/2=\pm 1/2$, a generic operator now has the following form:
$\sum c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\lambda_{\mu}.$
@@
$\sum c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\lambda_{\mu}$ $=\sum e^{i\pi(n_{1}-n_{2}+\mu)/2}c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\lambda_{\mu},$
$\sum c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\lambda_{\mu}$ $=\sum(-1)^{n_{2}+(\mu-1)/2}c^{\mu n_{1}n_{2}}\partial^{n_{2}}_{+\dot{+}}\partial^{n_{1}}_{-\dot{+}}\lambda_{-\mu}.$
The first constraint shows that when $\mu=1$ we have
$n_{1}-n_{2}\equiv 3\pmod{4},$
and the second constraint shows that The $\mu=-1$ coefficients $c^{-1,n_{1}n_{2}}$ are completely determined from the $\mu=1$ coefficients $c^{1,n_{1}n_{2}}$. From Table 3 we find that the fields $\lambda$ carry a $t^{3}$ factor, and therefore, the contribution of the $\lambda_{\pm}$ fields to the index is
$-3t^{3}\sum_{\tiny n_{1}-n_{2}\equiv 3\!\!\!\pmod{4}}t^{3n_{1}+3n_{2}}=-\frac{3t^{6}(1+t^{3})}{(1-t^{6})(1-t^{12})},$ (67)
We must also subtract from (67) the contribution of all operators made by acting with any number of $\partial_{\pm\dot{+}}$ on ++˙$\partial_{+\dot{+}}\lambda^{1}_{-}-\partial_{-\dot{+}}\lambda^{1}_{+}=\partial_{\alpha\dot{+}}\lambda^{1\alpha}$. Those operators are identically zero by the Dirac equation. Since $\partial_{\alpha\dot{+}}\lambda^{1\alpha}$ is invariant under $\Gamma$, we can account for this correction by simply subtracting $-t^{6}$ from $I_{\Phi}$, according to the charges of $\partial_{\alpha\dot{+}}\lambda^{1\alpha}$ in Table 3. Adding (66), with $I_{\Phi}$ replaced by $I_{\Phi}+t^{6}$, and (67), we find that the vector multiplet contribution to the single letter index for holonomies 1-4 is
$I^{vec}_{1}$ $=\frac{3(1-t^{6}+t^{12})(t^{2}-t^{4}+2t^{6})-3(1+t^{3})t^{6}}{(1-t^{6})(1-t^{12})}.$
Next, we compute $I^{vec}_{5}$, which is the vector multiplet single letter index for the $5^{th}$ holonomy of Table 4. This case is harder because the $5^{th}$ holonomy is nonabelian. The transformations of the adjoint field components $\Phi_{p}$ ($p=-1,0,1$) under the $SU(2)$ gauge transformations $g_{5}(r)$ and $g_{5}(s)$ are given in (50)-(51), which can be written more concisely as
$g_{5}(r)\Phi_{p}g_{5}(r)^{-1}=(-1)^{p}\Phi_{p}\,,\qquad g_{5}(s)\Phi_{p}g_{5}(s)^{-1}=-\Phi_{-p}\,.$
As before, we perform the computation for fields with spin projection $j_{1}=0$ and $j_{1}=\pm 1/2$ separately. A generic operator that corresponds to a spin $j_{1}=0$ field has the form:
$\sum c^{pn_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{p}.$
@@
$\sum e^{i\pi(n_{1}-n_{2})/2}c^{pn_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{p}$ $=\sum(-1)^{p}c^{pn_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{p},$
$\sum(-1)^{n_{2}}c^{pn_{1}n_{2}}\partial^{n_{2}}_{+\dot{+}}\partial^{n_{1}}_{-\dot{+}}\Phi_{p}$ $=-\sum c^{pn_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{-p}.$
@@
$e^{i\pi(n_{1}-n_{2})/2}c^{pn_{1}n_{2}}=(-1)^{p}c^{pn_{1}n_{2}},\qquad(-1)^{n_{1}}c^{-p,n_{2}n_{1}}=-c^{pn_{1}n_{2}}.$ (68)
The second constraint expresses $c^{(-1)n_{1}n_{2}}$ in terms of $c^{1n_{2}n_{1}}$, so we only need to consider $p=0,1$. We now perform the counting for $p=0$ and $p=1$ separately, and then add up the two contributions.
  • For $p=0$, the leftmost constraint of (68) requires
    $n_{1}-n_{2}\equiv 0\pmod{4},$
    and the second constraint expresses $c^{pn_{1}n_{2}}$ for $n_{1}<n_{2}$ in terms of $c^{pn_{2}n_{1}}$, and restricts $n_{1}$ to be odd when $n_{1}=n_{2}$. Thus, the number of independent parameters $c^{pn_{1}n_{2}}$ is the number of pairs $(n_{1},n_{2})$ with $n_{1}>n_{2}$ and $n_{1}\equiv n_{2}\pmod{4}$ plus the number of odd $n_{1}$’s. The contribution of those to the sum is
    $F^{0}=\frac{t^{6}}{(1-t^{6})(1-t^{12})}\left[I_{\Phi}(t)-(-t^{6})\right].$ (69)
    Here, we have already subtracted the term $(-t^{6})$ corresponding to the contribution of the Dirac equation $\partial_{\alpha\dot{+}}\lambda^{1\alpha}=0$, in order to not over count operators constructed from derivatives of terms with $j_{1}=\pm\frac{1}{2}$ (which will be counted later on).
  • For $p=1$, the leftmost constraint of (68) shows that
    $n_{1}-n_{2}\equiv 2\pmod{4}.$
    The contribution to the sum is therefore
    $F^{+}=\frac{2t^{6}}{(1-t^{6})(1-t^{12})}\left[I_{\Phi}(t)-(-t^{6})\right]=2F^{0}.$ (70)